# Durian: Dual Reference-guided Portrait Animation with Attribute Transferthanks: Project page: https://hyunsoocha.github.io/durian

### v1

@@ 3.2 Model Architecture: Dual ReferenceNet
where $\bar{\mathbf{F}}$ is the input feature map, $\bm{\phi}$ is the conditioning embedding, and $\mathbf{W}^{\prime}_{Q},\mathbf{W}^{\prime}_{K},\mathbf{W}^{\prime}_{V}$ are learned linear projections. Let $\bar{\mathbf{F}}_{\mathrm{attr}}^{l}$ and $\bar{\mathbf{F}}_{\mathrm{port}}^{l}$ denote the feature maps after self-attention in the $l$-th block of ARNet and PRNet, and let $\bar{\mathbf{F}}_{t}^{l}$ denote the feature map of DNet after spatial attention. The cross-attention updates are given by:
$\tilde{\mathbf{F}}_{\{\mathrm{attr,port}\}}^{l}=\mathrm{CA}(\bar{\mathbf{F}}_{\{\mathrm{attr,port}\}}^{l},\,\bm{\phi}_{\{\mathrm{attr,port}\}}),\tilde{\mathbf{F}}_{t}^{\tau,l}=\mathrm{CA}(\bar{\mathbf{F}}_{t}^{\tau,l},\,\bm{\phi}_{\mathrm{port}}),$ (9)
where $\tilde{\mathbf{F}}_{\mathrm{attr}}^{l}$, $\tilde{\mathbf{F}}_{\mathrm{port}}^{l}$, and $\tilde{\mathbf{F}}_{t}^{\tau,l}$ are the feature maps after cross-attention in ARNet, PRNet, and DNet, respectively.
@@
In addition to spatial and semantic conditioning, our model incorporates temporal awareness to generate coherent portrait animations. To model inter-frame consistency, we insert temporal self-attention into each block of the denoising U-Net (Hu, 2024; Zhu et al., 2024), enabling the model to capture motion dynamics across frames. To control pose and expression, we use a sequence of facial keypoints $\mathbf{K}={\bm{k}_{\tau}}_{\tau=1}^{F}$ extracted by Sapiens (Khirodkar et al., 2024). Each keypoint image $\bm{k}_{\tau}$ is encoded into a spatial feature map $\mathbf{F}_{\mathrm{kpt}}^{\tau}$ via a pose encoder and combined with the noisy latent $\bm{z}_{t}^{(\tau)}$ following (Zhu et al., 2024). For each frame $\tau$, our DNet $\epsilon_{\theta}$ predicts the added noise $\hat{\epsilon}^{(\tau)}_{t}$ from the noisy latent $\bm{z}_{t}^{(\tau)}$ at timestep $t$, using the reference features, semantic embeddings, and keypoint features:
@@
The predicted noise is used to recover the denoised latent $\bm{z}_{0}^{(\tau)}$, then decoded by the VAE decoder $\mathcal{D}$ to produce the final video frame:
$\mathbf{I}_{\tau}=\mathcal{D}(\bm{z}_{0}^{(\tau)}),\quad\text{for }\tau=1,\dots,F.$ (11)
This frame-wise prediction process, guided by temporal keypoints and reference features, enables the model to generate temporally coherent and pose-aligned portrait animations.
3.3 Training Strategy.
@@
To effectively train our model, we adopt a two-stage training scheme following previous approaches (Hu, 2024; Zhu et al., 2024). In the first stage, we optimize the entire model except the temporal attention layers using single-frame inputs. We define the per-frame conditioning bundle as $\mathcal{C}:=\left(\mathcal{F}_{\mathrm{attr}},\mathcal{F}_{\mathrm{port}},\bm{\phi}_{\mathrm{attr}},\bm{\phi}_{\mathrm{port}}\right),$ where $\mathcal{F}_{\mathrm{port}},\mathcal{F}_{\mathrm{attr}}$ are the multi-scale spatial features from PRNet and ARNet and $\bm{\phi}_{\mathrm{port}},\bm{\phi}_{\mathrm{attr}}$ are the semantic embeddings. Then, the training objective is the standard denoising diffusion loss:
@@
where $\bm{z}_{t}$ is the noised latent at diffusion timestep $t$, $\epsilon$ is the sampled noise, and $\mathbf{k}_{\tau}$ is the facial keypoint image at frame $\tau$. In the second stage, we freeze all modules except the temporal attention layers and train them using multi-frame inputs. The temporal objective considers a sequence of noised latents and corresponding keypoints:
@@
To simulate diverse spatial extents of facial attributes during training, we introduce an attribute-aware mask expansion strategy, illustrated in the top right of Fig.˜2. Given a training frame $\mathbf{I}$, we begin by selecting a target attribute (e.g., hair, eyeglasses, beard) and obtaining its binary segmentation mask $\mathbf{M}_{\mathrm{attr}}$ via Sapiens (Khirodkar et al., 2024). To simulate variation in the shape and coverage of this attribute, we generate an image $\mathbf{I}_{\mathrm{gen}}$ using SDXL (Podell et al., 2023) with ControlNet (Zhang et al., 2023), conditioned on the facial keypoints of $\mathbf{I}$ and a text prompt describing a modified attribute appearance (e.g., “long wavy hair”). We then extract a new attribute mask $\mathbf{M}_{\mathrm{gen}}$ using Sapiens. We define the initial candidate region as $\mathbf{M}_{\mathrm{port}}^{\mathrm{init}}:=\mathbf{M}_{\mathrm{attr}},$ and compute the expanded mask as the union of the initial and generated masks:
$\mathbf{M}_{\mathrm{port}}^{\mathrm{train}}=\mathbf{M}_{\mathrm{port}}^{\mathrm{init}}\cup\mathbf{M}_{\mathrm{gen}}.$ (14)
We then construct the two masked inputs used by the model:
$\tilde{\mathbf{I}}_{\mathrm{attr}}=\mathbf{I}\odot\mathbf{M}_{\mathrm{attr}},\quad\tilde{\mathbf{I}}_{\mathrm{port}}=\mathbf{I}\odot(1-\mathbf{M}_{\mathrm{port}}^{\mathrm{train}}),$ (15)
where $\odot$ denotes element-wise multiplication. The mask $\mathbf{M}_{\mathrm{port}}^{\mathrm{train}}$ defines the region into which the attribute will be inserted during generation, while $\mathbf{M}_{\mathrm{attr}}$ localizes the original attribute region. This expansion process is attribute-aware in that it preserves the intended attribute category while diversifying its spatial extent. Unlike HairFusion (Chung et al., 2025), which expands masks using fixed heuristics specific to hair, our approach generalizes across multiple facial attributes and enables the model to learn spatially flexible yet semantically grounded transfer patterns.
@@
To overcome the limited diversity inherent in self-reconstruction setups, we introduce an augmentation pipeline that improves robustness to pose, alignment, and appearance variations in attribute–portrait pairs. Our approach introduces controlled perturbations to both the attribute-only image and the attribute-masked portrait image to simulate realistic spatial and photometric variability. Specifically, we apply random affine transformations—including translation, scaling, and rotation—to induce spatial misalignment. To inpaint regions exposed by such transformations, we employ the FLUX outpainting model (Labs, 2024) for semantically consistent content completion. In addition, we apply color jittering operations such as random perturbation of tone, contrast, saturation, and hue to account for distribution shifts across image pairs. This augmentation strategy exposes the model to a wider range of training configurations, leading to more robust attribute transfer and animation under real-world variations.
Figure 3: Aligned Attribute Mask Estimation. To enhance spatial alignment between the attribute and the portrait input, we generate an aligned attribute mask by reenacting a 3D avatar reconstructed from the attribute image.
@@ 3.4 Inference Framework and Extensions
At inference time, we automatically construct the attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}$ and the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}$ by applying segmentation masks predicted by Sapiens (Khirodkar et al., 2024) to the attribute image $\mathbf{I}_{\mathrm{attr}}$ and the portrait image $\mathbf{I}_{\mathrm{port}}$, respectively. To improve spatial alignment between the attribute and portrait inputs, we generate an aligned attribute mask by animating a 3D avatar of the attribute image using the pose of the portrait. Concretely, we leverage GAGAvatar (Chu & Harada, 2024), a single-image-to-3D avatar pipeline, to reconstruct an avatar from $\mathbf{I}_{\mathrm{attr}}$, and animate it using the FLAME (Li et al., 2017) parameters $(\bm{\beta},\bm{\theta},\bm{\psi})$ of $\mathbf{I}_{\mathrm{port}}$ estimated via EMOCA (Daněček et al., 2022). From the resulting pose-aligned image $\mathbf{I}_{\mathrm{attr}}^{\mathrm{align}}$, we extract a new attribute mask $\mathbf{M}_{\mathrm{attr}}^{\mathrm{align}}$ using Sapiens. We then compute the union of this mask with the initial attribute mask extracted from the portrait image, denoted $\mathbf{M}_{\mathrm{port}}^{\mathrm{init}}$, to obtain the final candidate transfer region:
$\mathbf{M}_{\mathrm{port}}^{\mathrm{infer}}=\mathbf{M}_{\mathrm{port}}^{\mathrm{init}}\cup\mathbf{M}_{\mathrm{attr}}^{\mathrm{align}}.$ (16)
This updated mask is used to construct the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}=\mathbf{I}_{\mathrm{port}}\odot(1-\mathbf{M}_{\mathrm{port}}^{\mathrm{infer}})$. Fig.˜3 describes the overall mask estimation process. Then, we extract spatial features $\mathcal{F}_{\mathrm{attr}},\mathcal{F}_{\mathrm{port}}$ and semantic embeddings $\bm{\phi}_{\mathrm{attr}},\bm{\phi}_{\mathrm{port}}$ from the two masked reference images. Conditioned on these features and a keypoint sequence, DNet synthesizes an animation of the target identity with the desired attribute transferred by iterative denoising following Eq.˜10.
@@
Our model supports zero-shot composition of multiple attributes by generalizing the spatial attention formulation in Eq.˜6. Instead of using a single attribute feature, we concatenate multiple attribute feature maps along the width dimension:
@@
where each $\mathbf{F}_{\mathrm{attr}}^{l,k}$ denotes the feature map extracted from the $k$-th attribute-only image using the Attribute ReferenceNet. To construct the final attribute-masked portrait in this setting, we also generalize the mask fusion process by taking the union of all aligned attribute masks:
@@
Table 1: Quantitative Comparison. We compare our method with recent approaches that (1) generate portrait images from hair images, and (2) animate the synthesized portrait image. Our method outperforms both state-of-the-art baselines across all metrics.
@@
PbE LivePortrait 0.1059 16.14 0.5641 0.2859 40.63
X-Portrait 0.1180 15.33 0.5270 0.2978 59.20
MegActor-$\sum$ 0.1268 14.82 0.4840 0.3157 62.77
HairFusion LivePortrait 0.1438 13.76 0.4801 0.3792 46.24
X-Portrait 0.1511 13.30 0.4334 0.3733 59.02
MegActor-$\sum$ 0.1650 12.75 0.4138 0.4015 65.59
StableHair LivePortrait 0.1122 15.84 0.5491 0.3041 43.74
X-Portrait 0.1229 15.04 0.5114 0.3117 53.36
MegActor-$\sum$ 0.1301 14.62 0.4706 0.3347 63.47
TriplaneEdit LivePortrait 0.1023 16.52 0.5511 0.2924 57.86
X-Portrait 0.1051 16.05 0.5401 0.2760 60.25
MegActor-$\sum$ 0.1248 15.10 0.4828 0.3293 70.41
Ours 0.0744 18.83 0.6527 0.1565 38.00
Our model enables zero-shot interpolation between two attributes of the same category (e.g., hairstyle A and B) without additional training (Zhang et al., 2024a). Given two attribute-only images, we extract spatially attended features $\bar{\mathbf{F}}_{t}^{\tau,l,1}$ and $\bar{\mathbf{F}}_{t}^{\tau,l,2}$ using our spatial attention, and interpolate them as follows:
$\bar{\mathbf{F}}_{t}^{\tau,l}=(1-\alpha)\,\bar{\mathbf{F}}_{t}^{\tau,l,1}+\alpha\,\bar{\mathbf{F}}_{t}^{\tau,l,2}.$ (19)
, where $\alpha\in[0,1]$ controls the interpolation ratio. The interpolated feature $\bar{\mathbf{F}}_{t}^{\tau,l}$ is then passed to DNet for generation. This enables smooth and semantically consistent transitions between attributes.
4 Experiments
Table 2: Ablation Study. We evaluate the contributions of the key components and training strategies of our model. Bold text indicates the first rank, and underlined text denotes the second rank.

### v2

@@ 3.2 Model Architecture: Dual ReferenceNet
where $\bar{\mathbf{F}}$ is the input feature map, $\bm{\phi}$ is the conditioning embedding, and $\mathbf{W}^{\prime}_{Q},\mathbf{W}^{\prime}_{K},\mathbf{W}^{\prime}_{V}$ are learned linear projections. Let $\bar{\mathbf{F}}_{\mathrm{attr}}^{l}$ and $\bar{\mathbf{F}}_{\mathrm{port}}^{l}$ be the self-attended features of the $l$-th block in ARNet and PRNet, and $\bar{\mathbf{F}}_{t}^{l}$ the spatially attended feature of DNet. Then, the cross-attention updates are given by:
$\tilde{\mathbf{F}}_{\{\mathrm{attr,port}\}}^{l}=\mathrm{CA}(\bar{\mathbf{F}}_{\{\mathrm{attr,port}\}}^{l},\,\bm{\phi}_{\{\mathrm{attr,port}\}}),\quad\tilde{\mathbf{F}}_{t}^{\tau,l}=\mathrm{CA}(\bar{\mathbf{F}}_{t}^{\tau,l},\,\bm{\phi}_{\mathrm{port}}),$ (7)
where $\tilde{\mathbf{F}}_{\mathrm{attr}}^{l}$, $\tilde{\mathbf{F}}_{\mathrm{port}}^{l}$, and $\tilde{\mathbf{F}}_{t}^{\tau,l}$ are the feature maps after cross-attention in ARNet, PRNet, and DNet.
@@
Our model incorporates temporal awareness to generate coherent portrait animations by inserting temporal self-attention into each U-Net block, following Hu (2024); Zhu et al. (2024). To control pose and expression, we use a sequence of facial keypoints $\mathbf{K}=\{\bm{k}_{\tau}\}_{\tau=1}^{F}$ extracted by Sapiens (Khirodkar et al., 2024). Each keypoint image $\bm{k}_{\tau}$ is encoded into a spatial feature map $\mathbf{F}_{\mathrm{kpt}}^{\tau}$ via a pose encoder and combined with the noisy latent $\bm{z}_{t}^{(\tau)}$ following Zhu et al. (2024). For each frame $\tau$, DNet $\epsilon_{\theta}$ predicts the added noise $\hat{\epsilon}^{(\tau)}_{t}$ from the noisy latent $\bm{z}_{t}^{(\tau)}$ at timestep $t$, using the reference features, semantic embeddings, and keypoint features:
@@
The predicted noise is used to recover the denoised latent $\bm{z}_{0}^{(\tau)}$, then decoded by the VAE decoder $\mathcal{D}$ to produce the final video frame as $\mathbf{I}_{\tau}=\mathcal{D}(\bm{z}_{0}^{(\tau)})$ for $\tau=1,\dots,F$.
3.3 Training Strategy
@@
To effectively train our model, we adopt a two-stage training scheme following the previous approaches (Hu, 2024; Zhu et al., 2024). In the first stage, we optimize the entire model except the temporal attention layers, treating each video frame as an independent training sample. We define the per-frame conditioning bundle as $\mathcal{C}:=\left(\mathcal{F}_{\mathrm{attr}},\mathcal{F}_{\mathrm{port}},\bm{\phi}_{\mathrm{attr}},\bm{\phi}_{\mathrm{port}}\right),$ where $\mathcal{F}_{\mathrm{port}},\mathcal{F}_{\mathrm{attr}}$ are the multi-scale spatial features from PRNet and ARNet and $\bm{\phi}_{\mathrm{port}},\bm{\phi}_{\mathrm{attr}}$ are the semantic embeddings. Then, the training objective is the standard denoising diffusion loss:
@@
where $\bm{z}_{t}$ is the noised latent at diffusion timestep $t$, $\epsilon$ is the sampled noise, and $\mathbf{F}_{\mathrm{kpt}}$ is the feature map of the corresponding facial keypoint image. In the second stage, we freeze all modules except the temporal attention layers and train them using multi-frame inputs. The temporal objective considers a sequence of noised latents and corresponding keypoints:
@@
To expose the model to diverse spatial extents of facial attributes during training, we introduce an attribute-aware mask expansion strategy, illustrated in the top right of Fig.˜2. Given a training frame $\mathbf{I}$, we first select a target attribute (e.g., hair, eyeglasses, beard) and obtain its binary mask $\mathbf{M}_{\mathrm{attr}}$ using Sapiens (Khirodkar et al., 2024). To simulate variation in the shape and coverage of this attribute, we generate a modified image $\mathbf{I}_{\mathrm{gen}}$ with SDXL (Podell et al., 2023) and ControlNet (Zhang et al., 2023), conditioned on the facial keypoints of $\mathbf{I}$ and a text prompt describing an altered appearance (e.g., “long wavy hair”). A new mask $\mathbf{M}_{\mathrm{gen}}$ is then extracted from $\mathbf{I}_{\mathrm{gen}}$ using Sapiens. The final training mask is computed as the union of the original and generated masks, and the two masked inputs are constructed as:
$\mathbf{M}_{\mathrm{port}}^{\mathrm{train}}=\mathbf{M}_{\mathrm{attr}}\cup\mathbf{M}_{\mathrm{gen}},\quad\tilde{\mathbf{I}}_{\mathrm{attr}}=\mathbf{I}\odot\mathbf{M}_{\mathrm{attr}},\quad\tilde{\mathbf{I}}_{\mathrm{port}}=\mathbf{I}\odot(1-\mathbf{M}_{\mathrm{port}}^{\mathrm{train}}),$ (11)
where $\odot$ denotes element-wise multiplication. Here, $\mathbf{M}_{\mathrm{attr}}$ localizes the original attribute region, while $\mathbf{M}_{\mathrm{port}}^{\mathrm{train}}$ defines the expanded region into which the attribute will be inserted during generation. This expansion process is attribute-aware as it preserves the intended attribute category while diversifying its spatial extent. Unlike HairFusion (Chung et al., 2025), which expands masks using fixed heuristics specific to hair, our approach generalizes across multiple facial attributes and enables the model to learn spatially flexible yet semantically grounded transfer patterns.
@@
To address the limited diversity of self-reconstruction setups, we introduce an augmentation pipeline that improves robustness to pose, alignment, and appearance variations in attribute–portrait pairs. We perturb both the attribute-only and masked portrait images to simulate realistic spatial and photometric variations. We apply random affine transformations (translation, scaling, rotation) to induce spatial misalignment, and use the FLUX outpainting model (Labs, 2024) to inpaint newly exposed regions. Additionally, color jittering on tone, contrast, saturation, and hue accounts for appearance variations. This strategy exposes the model to diverse configurations, enabling more robust attribute transfer and animation under real-world variations.
Figure 3: Aligned Attribute Mask Estimation. To improve attribute-portrait alignment, we estimate an aligned attribute mask via Face Aligner.
@@ 3.4 Inference Framework and Extensions
At inference time, our system takes as input a portrait image, an attribute image, and a keypoint sequence. We first construct two masked reference images: the attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}$ and the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}$, by applying segmentation masks predicted by Sapiens (Khirodkar et al., 2024) to the attribute image $\mathbf{I}_{\mathrm{attr}}$ and the portrait image $\mathbf{I}_{\mathrm{port}}$. To improve spatial alignment between the attribute and portrait inputs, we introduce a Face Aligner module, which repurposes a lightweight image-to-3D avatar model (Chu & Harada, 2024) solely for alignment. This module reconstructs a coarse 3D avatar from the attribute image and aligns its shape and pose to the portrait using FLAME (Li et al., 2017) parameters $(\bm{\beta},\bm{\theta},\bm{\psi})$ estimated by EMOCA (Daněček et al., 2022). From the resulting pose-aligned image $\mathbf{I}_{\mathrm{attr}}^{\mathrm{align}}$, we extract a refined attribute mask $\mathbf{M}_{\mathrm{attr}}^{\mathrm{align}}$ using Sapiens. This mask is then merged with the initial portrait mask $\mathbf{M}_{\mathrm{port}}^{\mathrm{init}}$ to define the final transferable region $\mathbf{M}_{\mathrm{port}}^{\mathrm{infer}}=\mathbf{M}_{\mathrm{port}}^{\mathrm{init}}\cup\mathbf{M}_{\mathrm{attr}}^{\mathrm{align}}$. The updated mask is applied to construct the final attribute-masked portrait image, $\tilde{\mathbf{I}}_{\mathrm{port}}=\mathbf{I}_{\mathrm{port}}\odot(1-\mathbf{M}_{\mathrm{port}}^{\mathrm{infer}})$, as illustrated in Fig.˜3. Finally, spatial features $\mathcal{F}_{\mathrm{attr}},\mathcal{F}_{\mathrm{port}}$ and semantic embeddings $\bm{\phi}_{\mathrm{attr}},\bm{\phi}_{\mathrm{port}}$ are extracted from the two masked reference images. Conditioned on these features and the keypoint sequence, DNet synthesizes a video of the target identity with the desired attribute through iterative denoising (Eq.˜8).
@@
Our model supports zero-shot composition of multiple attributes without additional training, by generalizing the spatial attention formulation in Eq.˜5. Instead of using a single attribute feature, we concatenate multiple attribute feature maps along the width dimension:
@@
where each $\mathbf{F}_{\mathrm{attr}}^{l,k}$ denotes the feature map extracted from the $k$-th attribute-only image using the ARNet. To construct the final attribute-masked portrait in this setting, we also generalize the mask fusion process by taking the union of all aligned attribute masks:
@@
Our model enables zero-shot interpolation between two attributes of the same category (e.g., hairstyle A and B) without fine-tuning (Zhang et al., 2024a; Cha et al., 2025). Given two attribute-only images, we extract spatially attended features $\bar{\mathbf{F}}_{t}^{\tau,l,1}$ and $\bar{\mathbf{F}}_{t}^{\tau,l,2}$ using our spatial attention, and interpolate them as follows:
$\bar{\mathbf{F}}_{t}^{\tau,l}=(1-\alpha)\,\bar{\mathbf{F}}_{t}^{\tau,l,1}+\alpha\,\bar{\mathbf{F}}_{t}^{\tau,l,2},$ (14)
where $\alpha\in[0,1]$ controls the interpolation ratio. The interpolated feature $\bar{\mathbf{F}}_{t}^{\tau,l}$ is then passed to DNet for generation. This enables smooth and semantically consistent transitions between attributes.
4 Experiments
Table 1: Quantitative Comparison. We compare our method with recent approaches that (1) synthesize portraits with transferred hairstyles, and (2) animate the synthesized portrait image.
Self-Attribute Transfer Cross-Attribute Transfer
@@ 4 Experiments
PbE LivePortrait 0.1059 16.14 0.5641 0.2859 40.63 0.8499 0.6407 0.5630 37.6462 3.3868
X-Portrait 0.1180 15.33 0.5270 0.2978 59.20 0.8393 0.5916 0.5458 36.7030 2.9008
MegActor-$\sum$ 0.1268 14.82 0.4840 0.3157 62.77 0.8535 0.6266 0.4863 38.2746 6.2743
HairFusion LivePortrait 0.1438 13.76 0.4801 0.3792 46.24 0.8741 0.6843 0.6502 30.5632 2.6719
X-Portrait 0.1511 13.30 0.4334 0.3733 59.02 0.8809 0.6914 0.6520 30.2570 4.9184
MegActor-$\sum$ 0.1650 12.75 0.4138 0.4015 65.59 0.8736 0.6708 0.6044 30.9702 5.3037
StableHair LivePortrait 0.1122 15.84 0.5491 0.3041 43.74 0.8831 0.7051 0.6564 29.5014 3.9495
X-Portrait 0.1229 15.04 0.5114 0.3117 53.36 0.8895 0.7239 0.6443 28.2627 1.5718
MegActor-$\sum$ 0.1301 14.62 0.4706 0.3347 63.47 0.8848 0.7271 0.6130 30.4087 1.4672
TriplaneEdit LivePortrait 0.1023 16.52 0.5511 0.2924 57.86 0.8540 0.6163 0.2776 32.5660 8.9103
X-Portrait 0.1051 16.05 0.5401 0.2760 60.25 0.8366 0.6216 0.2944 30.6319 2.9315
MegActor-$\sum$ 0.1248 15.10 0.4828 0.3293 70.41 0.8210 0.5674 0.2770 32.5679 2.8542
Ours 0.0744 18.83 0.6527 0.1565 38.00 0.9043 0.7801 0.7098 27.1547 2.4052
Figure 4: Qualitative Comparison for Cross-Attribute Transfer. We compare our method and the baselines that combine X-Portrait (Xie et al., 2024) with StableHair (Zhang et al., 2025) in cross-identity transfer setup. We provide more results in our Supp. Mat.
Experimental setup.
To address the lack of ground-truth data for cross-identity attribute transfer, we design two evaluation settings: self-attribute transfer and cross-attribute transfer. In self-attribute transfer, a single video is split into a portrait and an attribute image from different frames of the same identity, and the model reconstructs the original video. While useful for controlled evaluation, this provides only a pseudo ground-truth and mainly reflects reconstruction ability rather than the full complexity of cross-identity transfer. In cross-attribute transfer, the portrait and attribute images come from different individuals. Without exact ground-truth, this setting instead evaluates semantic consistency, identity preservation, and temporal realism. Together, the two settings offer a comprehensive evaluation of both low-level fidelity and high-level transfer quality.
Dataset.
We train our model on CelebV-Text (Yu et al., 2023), VFHQ (Xie et al., 2022), and Nersemble (Kirschstein et al., 2023), totaling 2,747 videos. For evaluation, we sample 200 videos for self-attribute transfer and 50 videos for cross-attribute transfer from CelebV-Text and VFHQ, ensuring diverse and unseen identities, head poses, and expressions. The masks for the portrait and attribute frames are generated following the procedure used in each compared method.
Metrics.
For self-attribute transfer, we evaluate reconstruction fidelity using $\text{L}_{1}$, PSNR, SSIM, and LPIPS, and perceptual quality with FID (Parmar et al., 2022). For cross-attribute transfer, we measure attribute transfer quality with CLIP-I (Radford et al., 2021; Hessel et al., 2021) and DINO (Caron et al., 2021), identity preservation with ArcFace (Deng et al., 2019), and temporal realism with VFID (Fang et al., 2024) using I3D (Carreira & Zisserman, 2017) and ResNeXt (Hara et al., 2018).
4.1 Comparison
Baselines.
As no prior work directly tackles portrait animation with attribute transfer from in-the-wild reference images, we construct two-stage baselines by combining image-level attribute transfer with video animation methods, resulting in 12 model combinations. For attribute transfer (stage 1), we consider: Paint-by-Example (PbE) (Yang et al., 2023), a mask-conditioned diffusion method for reference image insertion; HairFusion (Chung et al., 2025) and StableHair (Zhang et al., 2025), diffusion-based models for hairstyle transfer with and without masks; and TriplaneEdit (Bilecen et al., 2024), a 3D-aware GAN-based face editor. For portrait animation (stage 2), we use: LivePortrait (Guo et al., 2024), X-Portrait (Xie et al., 2024), and MegActor-$\sum$ (Yang et al., 2025).
Results.
As shown in Table˜1, our method consistently outperforms all baseline combinations across both fidelity and perceptual quality metrics in self-attribute transfer. Fig.˜4 presents a qualitative comparison against baselines using LivePortrait (Guo et al., 2024) as the animation module (stage 2). Our method generates coherent and realistic hairstyle animations that preserve the identity and maintain consistency in spatial extent, shape, and fine details across frames. Please refer to our Supp. Mat. for additional qualitative comparisons with other baseline combinations.
4.2 Ablation Study
Table 2: Ablation Study. Bold indicates the best, underline the second.